OpenStudy (anonymous):

A very small unit of length is the femtometer (fm). One fm= 10^-15 meters and 2.54 cm = 1 inch. One cubic inch is how many cubic femtometers? (using dimensional analysis)

OpenStudy (nechirwan):

1 in^3 = (2.54 cm)^3 = (2.54 x 10^-2 m)^3 = 16.387064 x 10^-6 m = 16.387064 x 10^-6 m But 1 fm = 10^-15 m So 1 m = 1 x 10^15 fm Therefore converting a cubic inch to femtometres, 1 in^3 = (16.387064 x 10^-6 m) x (10^15 fm per m) =1.6387064 x 10^10 fm